OpenStudy (anonymous):

What is the surface area of this right rectangular prism? *Will medal and fan*

OpenStudy (anonymous):

Please explain how you get the area of this.

OpenStudy (sweetburger):

take the area of each side and add it together

OpenStudy (anonymous):

8^2

OpenStudy (anonymous):

8m2

OpenStudy (anonymous):

I didn't come here for the answer exactly, I actually want to know how to solve this with the formula.

OpenStudy (anonymous):

You want the formula or a deeply explanation of how to do it.

OpenStudy (sweetburger):

Wait the formula you provided was wrong...

OpenStudy (anonymous):

Find the area of adjacent sides (Width*Height)*2 sides. Find the area of ends (Length*Width)*2 ends. Add the three areas together to find the surface area. Example: The surface area of a rectangular prism 5 cm long, 3 cm. wide and 2 cm. high = 5*2*2 + 3*2*2 + 5*3*2 = 20 + 12 + 30 = 62 cm2.

OpenStudy (sweetburger):

it should be 2(wl+hl+hw)